OpenStudy (anonymous):

Write an equation in point slope form of the line that passes through (5, 0), and whose slope is -2.

OpenStudy (waleed_imtiaz):

U know point-slope form is y-y1=m(x-x1) so here, x1=5 and y1=0 and slope=-2 so put the values..... y-0=-2(x-5) so, y= -2x+10

OpenStudy (waleed_imtiaz):

or u can write it as 2x+y-10=0